A ball moving with speed ν hits another identical ball at rest, The two balls stick together after collision. If specific heat of the material of the balls is S, the temperature rise resulting from the collision is

Grade:12th pass

1 Answers

Sandeep Bhamoo
5 years ago
By cons. of momentum final vel.=v/2 And by energy cons. Initial kinetic energy= final kinetic energy + heat putting , mv²/2=mv²/4 + 2ms(∆T). ∆T= v²÷(8s). Please do approve my answer if correct